OpenStudy (souvik):

an electromagnetic wave travels in xy plane making an angle theta with x axis.what is the equation of the wave?

OpenStudy (vincent-lyon.fr):

Do you need real or complex expression? Do you need both E and B fields?

OpenStudy (souvik):

both E and B

OpenStudy (vincent-lyon.fr):

Well, the most interesting part is the phase of the wave. Do you have an idea what it looks like?

OpenStudy (vincent-lyon.fr):

If the wave had been travelling along the x-axis, its phase would have been: \(\omega t-kx\) with \(k=\omega /c\) Do you agree?

OpenStudy (souvik):

yes

OpenStudy (vincent-lyon.fr):

Now this part \(kx\) is actually the dot product of wave vector \(\vec k\) and position \(\vec r=\vec{OM}\) Hence: \(\omega t-\vec k .\vec r\) is the general expression.

OpenStudy (souvik):

okay..

OpenStudy (vincent-lyon.fr):

So now... can you make up the right expression using your angle \(\theta\) ?

OpenStudy (souvik):

k vector is directed to positive x direction..and r vector makes an angle theta with X axis..so the dot product is krcos(theta)?

OpenStudy (vincent-lyon.fr):

No, it is the other way round: \(\vec k\) is pointing at \(\theta\) with respect to the x-axis whereas \(\vec r\) is any point (x,y,z) in space. 

OpenStudy (vincent-lyon.fr):

Remember \(\vec k\) represents the direction of the wave, like a ray of light does.

OpenStudy (souvik):

then the phase is k*x*cos(theta)+k*y*sin(theta)-wt..............?

OpenStudy (vincent-lyon.fr):

Perfect! You can leave k as it is or replace it with \(\omega /c\) if the wave is travelling in a vacuum.

OpenStudy (souvik):

ok...thank u...you are great sir.....please tell me a good modern physics book...

OpenStudy (vincent-lyon.fr):

Sorry, I have no idea! I do not teach in English-speaking countries, so I do not know which books are good or not. What level (what age) are you referring to? I know one or two books for Relativity, but that's all I can do.

OpenStudy (souvik):

college level...i just passed high school...where are you from?

OpenStudy (vincent-lyon.fr):

I am from France. What about you?

OpenStudy (souvik):

i am from India..do you teach in college?

OpenStudy (vincent-lyon.fr):

Yes, I teach in 2nd year and I also teach teachers-in-training. This book seems good (particularly the first chapters on special relativity): Bernard Schutz, A first course in General Relativity, (Cambridge University Press).

OpenStudy (souvik):

thank you again sir....:)

OpenStudy (vincent-lyon.fr):

yw :))